Hillside issues $1,000,000 of 6%, 15-year bonds dated January 1, 2017, that pay interest semiannually on June 30 and DecemberReg 1 Reg 2A to 20 Reg 3 Red 4 Reg For each semiannual period, complete the table below to calculate the cash payment, straigReq 1 Req 2A to 20 Reqh Reg 4 Reg 5 Req3 Complete the below table to calculate the total bond interest expense to be recognizReq 1 Reg za to 2c Req 3 Reg 4 Reg 5 Prepare the first two years of an amortization table using the straight-line method ncesPrepare the journal entries to record the first two interest payments. View transaction list Journal entry worksheet 2 RecordJournal entry worksheet Record the second interest payment on December 31, 2017. Note: Enter debits before credits. General J
Answer #1
Journal entry
Date General Journal Debit Credit
1/1/2018 Cash 1,223,995
premium on bonds 223,995
bonds payable 1,000,000
2-a) par maturity value Annual rate / year semi annual cash payment
1,000,000 * 6% 6./12 30000
semi annual Straight line
2-b) bond price par value premium periods premium amortization
1,223,995 - 1,000,000 = 223,995 / 30 = 7467
2-c) Semi annual cash premium bond interest expense
payment amortization
30,000 - 7467 = 22,534
3) total bond interest expense over life of bonds
amount repaid
30 payments of 30,000 900000
par value ant maturity 1,000,000
total repaid 1900000
less amount borrowed 1,223,995
total bond interest expense. 676,005
(note bond interest expense may differ slightly due to rounding)
4) unamort Carrying
period premium value
1/1/2018 223,995 1,223,995
06/30/18 216,529 1,216,529
12/31/18 209,062 1,209,062
06/30/19 201,596 1,201,596
12/31/19 194,129 1,194,129
5)
Date General Journal Debit Credit
6/30/18 interest expense 22,534
premium on bonds payable 7,467
cash 30,000
31/12/2018
interest expense 22,534
premium on bonds payable 7,467
cash 30,000